OpenStudy (anonymous):

Can anyone help me with these? I've tried but dont get them at all. Any help ? y varies directly as x. If y = 12 when x = 8, find y when x = 6. y varies directly as x. y = 75 when x = 100. Find the constant of variation y varies directly as x^(2). If y is 150 when x is 5, find y when x is 6 y varies directly as x2. If y is 150 when x is 5, find y when x is 3. If y varies directly as x, and x = 4 and y = 8, find x when y = 2. y varies directly as x. If y = 12 when x = 3, find y when x = 6. y varies inversely as x. If y = 50 when x = - 3, find y when x = - 10. y varies inversely as x^(

OpenStudy (ivettef365):

use formula y=kx, to find k

OpenStudy (ivettef365):

y = 12 and x = 8, then you have 12=k8, therefore k=12/8, so k=3/2

OpenStudy (ivettef365):

now find y when x=6, you know k=3/2 so use formula again y=kx, so y=3/2(6) solve for y

OpenStudy (ivettef365):

k is the constant of variation